Accelerator Update for August 13, 2008 - August 15, 2008
Wednesday August 13
The day shift began with Operations monitoring the Tevatron (TeV) store (66.75E30), the Recycler (RR) stash (355.6E10), and the Antiproton Source (Pbar) stack (25.2E10), and with the BNB and NuMI experiments taking beam.
Operations terminated store 6356 at 11:23 AM.
Operations put the TeV into a shot setup at 11:43 AM.
Operations established store 6357 at 1:40 PM with an initial luminosity of 294.81E30.
At 1:48 PM, a controls expert installed a new Ecool frontend.
Beam was established to MTest at 6:04 PM.
Thursday August 14
The midnight shift began with Operations monitoring store 6357 (96.76E30) and the stash (227.5E10), with Pbar stacking (24.0E10), and with the BNB, NuMI and MTest experiments taking beam.
Operations terminated store 6357 at 4:17 AM.
Operations put the TeV into a shot setup at 4:26 AM.
Operations established store 6358 at 6:44 AM with an initial luminosity of 266.85E30.
At 6:47 AM, Linac, Booster, and Main Injector were prepared for access for maintenance and repair work.
Beam was re-established to the Linac Dump at 10:30 AM. Linac is conducting studies while waiting for Booster and Main Injector work to be completed.
At 11:46 AM there was a vacuum burst that started in the P150 line and got all the way to F0 before it was caught. This resulted in a loss of luminosity lifetime.
Beam was established to the Booster dump at 1:27 PM.
At 2:03 PM a water leak was found on the PBar lens transformer, experts decided to change out the lens, no PBar beam until this is completed.
Beam returned to the BooNEs at 2:38 PM.
Beam returned to NuMI at 4:15 PM.
Operations began turning MTest on at 8:00 PM, they were unable to run MT5E above 200 amps, an expert was called in to investigate.
At 11:00 PM, it was determined that one of the magnets in the MT5E string has an internal ground fault. The magnet will have to be replaced tomorrow morning. MTest will not take beam until this is repaired.
Friday August 15
The midnight shift began with Operations monitoring store 6358 (55.76E30) and with no stash. Pbar is down for the lens changeout, and with the BNB and NuMI taking beam.
Booster East Anode Power Supply tripped off at 12:06 AM, BRF3 was the culprit and has been removed from the circuit.
Booster beam returned at 12:22 AM.
The PBar Lens change was completed at 5:19 AM.
PBar stacking resumed at 5:50 AM.
The Plans for Friday and the weekend
The plans for today are to stack, stash, and store. Replace the bad MT5E magnet and get MTest back on line.
Comments and Suggestions
|last modified 08/18/2008 email Fermilab|
|Security, Privacy, Legal|